OpenStudy (kidthatbro8):

Which of the following is equal to square root of the cube root of 5? 5 to the power of 1 over 3 5 to the power of 1 over 6 5 to the power of 2 over 3 5 to the power of 3 over 2

OpenStudy (anonymous):

\[\sqrt{\sqrt[3]{5}}\] \[((5)^\frac{ 1 }{ 3 })^\frac{ 1 }{ 2 } = (5)^{\frac{ 1 }{ 3 }*\frac{ 1 }{2 }}\]

OpenStudy (kidthatbro8):

i'm not sure how to solve that...? @sangya21

OpenStudy (anonymous):

\[\frac{ 1 }{ 3 }* \frac{ 1 }{ 2 } = \frac{ 1 }{ 3*2 } = \frac{ 1 }{ 6 }\] \[5^\frac{ 1 }{ 6 }\]

OpenStudy (kidthatbro8):

oh, thank you !

OpenStudy (anonymous):

Anytime :)